the probability that a doctor diagnoses a particular illness is 0.7. given that the doctor makes an incorrect diagnoses, the probability that the patient enters a lawsuit is 0.9. what is the probability that the doctor makes an incorrect diagnoses and the patient sues?
Answer by stanbon(75887)

the probability that a doctor diagnoses a particular illness is 0.7. given that the doctor makes an incorrect diagnoses, the probability that the patient enters a lawsuit is 0.9. what is the probability that the doctor makes an incorrect diagnoses and the patient sues?
------
Note:: P(a AND b) = P(b)P(a|b)
-------------------
P(diags) = 0.7 ; P(wrong diags) = 0.3
---
P(lawsuit | wrong diag) = 0.9
------
P(wrong diag AND lawsuit) = 0.3*0.9 = 0.27
-------
Cheers,
Stan H.
----------------